The fundamental frequency(V0)of HCL molecule is 2890 cm.calculate the fundamental frequency (V0) for DCL. (Use atomic masses as H=1, D=2,Cl= 35)
1
Expert's answer
2018-05-17T08:35:43-0400
M(HCl) = m(H)*m(Cl)/m(H)+m(Cl)= 1*35/1+35=0.97
M(DCl) = 2*35/2+35=1.89
V0(HCl)/V0(DCl) = (M(DCl)/M(HCl))^1/2
V0(HCl)/V0(DCl) = 1,394
V0(DCl)=V0(HCl)/1.394=2890/1.394=2073.2 - frequency